# Fish Tracking, Counting, and Behaviour Analysis in Digital Aquaculture: A Comprehensive Review

### v1

@@ IV-C Fish behaviour analysis based on passive acoustic monitoring
Passive acoustic monitoring (PAM) has emerged as a non-invasive and increasingly accessible remote sensing technology for monitoring underwater environments [205, 206]. With approximately 1,000 out of the 35,000 known fish species confirmed to produce sounds underwater [207, 208], PAM offers a unique opportunity to analyze fish behaviour through the sounds they generate (The example of fish abnormal behaviour is shown in Fig. 8).
Fish can produce a series of sounds during feeding, and the frequency spectrum of these sounds can be used to analyze their feeding behaviour. For example, turbots generate feeding sounds that vary with food intake intensity, ranging from $15$ to $20$ dB in the frequency range 777710101010 kHz [209]. Similarly, feeding sounds produced by various fish species, such as rainbow trout (0.020.020.020.0225252525 kHz) [210], Japanese minnow (111110101010 kHz) [211], Atlantic horse mackerel (1.61.61.61.64444 kHz)[212], yellowtail (44446666 kHz) [213], have comparable frequency ranges.
The Fish feeding behaviour analysis based on audio was initially proposed by [4, 214], the audio signal is first transformed into log mel spectrograms and then fed into a CNN-based model for FFIA. Subsequent work [215, 216] have further demonstrated the feasibility of using audio as input for FFIA. Audio-based methods offer advantages such as energy efficiency and lower computational costs compared to vision-based methods [217, 218]. However, audio-based models have lower classification performance than video-based FFIA due to their inability to capture full visual information and sensitivity to environmental noise [219]. Moreover, rapidly swimming predatory fish, such as brown and rainbow trout, often combine forward swimming with feeding, accompanied by splashing sounds and strong tail patting [172]. The rapid pellet capture by these species superimposes feeding sounds, and pellet impacts pose a challenge in obtaining accurate feeding sound data.
@@
Figure 8: The audio spectrum of different fish abnormal behaviours [220].
@@ IV-D Fish behaviour analysis based on biosensor technology
The feeding behaviour of most fish leads to characteristic changes in acceleration that differ from their normal movement patterns [223]. These characteristic changes in acceleration can be effectively used to distinguish feeding behaviour patterns from other behaviour patterns [224]. For example, [225] used accelerometer tags to investigate the feeding behaviour of Atlantic cod (Gadus morhua) in the wild. The authors found that the accelerometer data could accurately identify feeding events and provide insights into the foraging ecology of this species. Similarly, a study by [226] used a combination of accelerometers and gyroscopes to analyze the feeding behaviour of captive yellowtail kingfish (Seriola lalandi). The authors demonstrated that the sensor data could be used to classify different types of feeding behaviour, such as biting, chewing, and swallowing, with high accuracy.
In addition to feeding behaviour, biosensors have been used to study other aspects of fish behaviour, such as swimming activity and energy expenditure. For instance, [227] used accelerometers to investigate the swimming behaviour and energy expenditure of wild Atlantic salmon (Salmo salar) while migrating to spawning grounds. The authors found that the accelerometer data provided valuable insights into the swimming performance and energy costs of this species in natural conditions.
@@
V Public dataset
@@
High-quality public datasets are crucial for developing and evaluating computer vision and deep learning methods for fish detection, tracking, and behaviour analysis. However, despite the growing popularity of deep learning, there are still relatively few public datasets specifically focused on underwater fish scenes. This scarcity has led many researchers to conduct their analyses and behavioural studies under ideal or controlled conditions. Table IV summarizes the available public fish datasets.
TABLE IV: Summary of the various fish datasets.
@@
$2704\times 1520$ 86452 Tracking [69]
@@
FISHTRAC 14 videos $1920\times 1080$ 3449
@@
VI Challenges and future perspectives
@@
1) Massively available datasets: The wide application of intelligent technology in aquaculture, especially the success of deep learning algorithms in image processing [15], has highlighted the need for large-labelled datasets. Although available datasets are gradually increasing, most are limited to identifying and detecting fish species. Open data on fish tracking, counting, and behaviour analysis is scarce. Passive acoustic monitoring (PAM) is also gaining popularity for underwater listening [239, 206] and public sound data of underwater fish (e.g., Fishsound) are emerging. However, the sample size of these datasets has not yet reached critical mass. In the future, developing an international platform for sharing images and acoustic data will be essential to promote sustainable aquaculture development.
2) Audio-visual multi-modal techniques: Fish tracking, counting, and behaviour analysis methods are limited to single modalities (acoustic or computer vision). However, the complex aquaculture environment leads to one-sided data that cannot fully capture all fish information [214]. Multimodal machine learning aims to establish models that process and associate data from multiple modalities. With the development of multimodal machine learning, crowd tracking and behaviour analysis based on audio-visual data have attracted extensive attention. Multimodal learning for fish is still in its infancy, but combining video data, sonar imaging data, and active acoustic data can better model fish tracking, counting, and behaviour quantification tasks, further improving the level of intelligence in aquaculture.
@@
3) On-device machine learning: Most current fish tracking, counting, and behavioural analysis models are performed in the cloud or on high-performance GPUs. However, many aquaculture tasks require real-time responses, such as fish feeding and abnormal behaviour detection. Cloud-based models may struggle to guarantee this real-time performance, and many devices in remote and harsh aquaculture environments may not have consistent internet connectivity. On-device models can greatly reduce exercise pressure and make devices more intelligent, providing users with a better experience. However, terminal devices have processing power, power consumption, cost, and volume limitations. Future developments could focus on reducing the complexity of computing and storage by optimizing neural network algorithms or compressing network models using techniques like knowledge distillation to enable direct running on device chips.
@@
4)Integration of fish tracking, counting, and behaviour analysis: Most research addresses fish tracking, counting, and behaviour analysis as separate tasks. However, these tasks are often interconnected in real-world aquaculture scenarios and must be performed continuously in the same environment. Developing a joint model that can handle all three tasks simultaneously would be more memory-efficient and suitable for practical applications in aquaculture. A joint model would leverage the shared features and information among the tasks, reducing redundancy and improving overall performance. For example, accurate fish tracking can provide valuable information for counting and behaviour analysis, while behaviour analysis can help identify and resolve tracking challenges such as occlusions and interactions between fish. Researchers can develop more comprehensive and efficient systems for monitoring and managing aquaculture farms by integrating these tasks into a single framework. This approach would also reduce the computational resources required, making it more feasible to deploy such systems in real-world settings. Future research should focus on developing novel architectures and training strategies that can effectively combine fish tracking, counting, and behaviour analysis tasks.
@@
VII Conclusion
This review provides a comprehensive analysis of the current state of digital technologies in aquaculture, including vision-based sensors, acoustic-based sensors, and biosensors, for fish tracking, counting, and behaviour analysis. These technologies offer valuable tools for optimizing production efficiency, fish welfare, and resource management in aquaculture. However, each technology has its own limitations, such as the sensitivity of vision-based sensors to environmental conditions, the high cost and complexity of acoustic-based sensors, and the potential invasiveness of biosensors. Despite the advancements made in these technologies, significant challenges remain, including the scarcity of comprehensive fish datasets, the lack of unified evaluation standards, and the need for more robust and adaptable systems that can handle the complexities of real-world aquaculture environments. To address these challenges and drive progress in the field, future research should focus on developing diverse and representative datasets, establishing standardized evaluation frameworks, and exploring the integration of multiple technologies to create more comprehensive and reliable monitoring systems. Furthermore, emerging technologies such as multimodal data fusion, deep learning, and edge computing present exciting opportunities for advancing digital aquaculture. By leveraging these technologies, researchers can develop more accurate, efficient, and practical solutions for fish tracking, counting, and behaviour analysis, ultimately contributing to the sustainable growth and development of the aquaculture industry.
VIII AUTHOR CONTRIBUTIONS

### v2

@@ IV-C Fish behaviour analysis based on passive acoustic monitoring
Passive acoustic monitoring (PAM) has emerged as a non-invasive and increasingly accessible remote sensing technology for monitoring underwater environments [254, 255]. With approximately 1,000 out of the 35,000 known fish species confirmed to produce sounds underwater [256, 257], PAM offers a unique opportunity to analyze fish behaviour through the sounds they generate. An audio example of fish abnormal behaviour is shown in Fig. 12.
Fish can produce a series of sounds during feeding, and the frequency spectrum of these sounds can be used to analyse their feeding behaviour. For example, turbots generate feeding sounds that vary with food intake intensity, ranging from $15$ to $20$ dB in the frequency range 777710101010 kHz [258]. Similarly, feeding sounds produced by various fish species, such as rainbow trout (0.020.020.020.0225252525 kHz) [259], Japanese minnow (111110101010 kHz) [260], Atlantic horse mackerel (1.61.61.61.64444 kHz)[261], yellowtail (44446666 kHz) [262], have comparable frequency ranges.
The Fish feeding behaviour analysis based on audio was initially proposed by [16, 263], where the audio signal is first transformed into log Mel spectrograms and then fed into a CNN-based model for FFIA. Subsequent work [264, 265] have further demonstrated the feasibility of using audio as input for FFIA. Audio-based methods offer advantages such as energy efficiency and lower computational costs compared to vision-based methods [266, 267]. However, audio-based models have lower classification performance than video-based FFIA due to their inability to capture full visual information and sensitivity to environmental noise [268]. Moreover, rapidly swimming predatory fish, such as brown and rainbow trout, often combine forward swimming with feeding, accompanied by splashing sounds and strong tail patting [217]. The rapid pellet capture by these species superimposes feeding sounds, and pellet impacts pose a challenge in obtaining accurate feeding sound data.
@@
Figure 12: The audio spectrum of different fish abnormal behaviours (This figure was reproduced from [269]).
@@ IV-D Fish behaviour analysis based on biosensor technology
The feeding behaviour of most fish leads to characteristic changes in acceleration that differ from their normal movement patterns [272]. These characteristic changes in acceleration can be effectively used to distinguish feeding behaviour patterns from other behaviour patterns [273]. For example, in [274], accelerometer tags were used to investigate the feeding behaviour of Atlantic cod (Gadus morhua) in the wild. The authors found that the accelerometer data could accurately identify feeding events and provide insights into the foraging ecology of this species. Similarly, a study by [275] used a combination of accelerometers and gyroscopes to analyse the feeding behaviour of captive yellowtail kingfish (Seriola lalandi). The authors demonstrated that the sensor data could be used to classify different types of feeding behaviour, such as biting, chewing, and swallowing, with high accuracy.
In addition to feeding behaviour, biosensors have been used to study other aspects of fish behaviour, such as swimming activity and energy expenditure. For instance, in [276], accelerometers were used to investigate the swimming behaviour and energy expenditure of wild Atlantic salmon (Salmo salar) while migrating to spawning grounds. The authors found that the accelerometer data provided valuable insights into the swimming performance and energy costs of this species in natural conditions.
@@
IV-E Comparative analysis and trends in fish school behaviour analysis
Fish school behaviour analysis in aquaculture focuses on understanding and monitoring fish activities, stress responses, and interactions within the aquaculture environment. This field employs various technologies, each with distinct capabilities and applications. Computer vision-based analysis excels in observing feeding patterns, detecting abnormal behaviours, and identifying signs of stress or disease [277]. Advanced algorithms can recognize specific behaviour patterns such as aggressive interactions, schooling dynamics, and responses to environmental changes. While being non-invasive, this method’s effectiveness can be limited by water clarity and lighting conditions [26]. Multi-object tracking extends computer vision capabilities by following individual fish within schools over time. This approach provides insights into social hierarchies, individual growth rates, and how behaviour patterns spread through a population. It requires sophisticated algorithms and high computational power but offers unparalleled detail in understanding group dynamics [11, 4].
Passive acoustic monitoring analyses sounds produced by fish to infer behaviour states and stress levels [254]. This method is particularly valuable for nocturnal species or in turbid environments where visual observation is challenging. It can detect feeding activity, spawning events, and even some types of distress calls, offering a unique perspective on fish behaviour [278]. Biosensor technology, including implantable tags and external sensors, provides detailed physiological data that can be correlated with behaviour [274]. These sensors can measure parameters like heart rate, muscle activity, and hormone levels, offering insights into stress responses, energy expenditure during different activities, and even feeding intensity [279]. However, the invasive nature of some biosensors raises welfare concerns and may itself affect behaviour. The choice of behaviour analysis method often depends on research goals, species-specific characteristics, and environmental conditions. For example, salmon farms might prioritize feeding behaviour analysis to optimize feed conversion ratios, while tilapia farms might focus more on aggression and territoriality behaviours.
Regional trends in adoption vary. Asia, with its diverse aquaculture species and systems, shows broad adoption across methods [280, 281]. Europe emphasizes non-invasive techniques due to strict welfare regulations [282]. North America leads in integrating multiple methods for comprehensive behaviour analysis Welfare considerations are increasingly central to method selection [283, 284]. There is a growing preference for non-invasive techniques that can provide detailed behavioural data without causing stress. This has spurred research into improving the resolution and accuracy of passive monitoring methods. The trend in fish behaviour analysis is moving towards holistic approaches that combine multiple methods to create a comprehensive picture of fish welfare and behaviour. For instance, combining computer vision with passive acoustics can provide round-the-clock monitoring that captures both visual and auditory behavioural cues [263]. Real-world applications have demonstrated the value of behaviour analysis in improving aquaculture management. Farms using advanced behaviour monitoring have reported earlier detection of disease outbreaks, more efficient feeding practices, and improved overall fish welfare, leading to better growth rates and product quality [285, 286]. As aquaculture intensifies and faces new challenges like climate change impacts, behaviour analysis will play an increasingly crucial role in maintaining fish health, optimizing production, and ensuring sustainable practices.
IV-F Real-world applications in fish behaviour analysis
Fish behaviour analysis in real-world applications has emerged as a powerful tool for monitoring aquaculture environmental conditions and fish welfare. By tracking and analysing changes in swimming patterns, spatial distribution, and movement characteristics, these systems can serve as early warning indicators for water quality issues and stress responses, enabling timely interventions in commercial aquaculture operations. Xu et al. [287] demonstrate how fish behaviour analysis can be used to monitor water quality by studying responses to ammonia nitrogen stress. Their system combines an improved YOLOv8 model (enhanced with multi-head self-attention and Wise Intersection over Union (Wise-IoU)) with dual-view cameras for 3D tracking. By integrating advanced tracking techniques including Kalman filtering and Kernelized Correlation Filters [288], the system accurately monitors multiple fish simultaneously. The study analysed behavioural patterns of different species (sturgeon, bass, and crucian) under various ammonia nitrogen stress levels, examining their trajectories, exercise volumes, spatial distribution, and movement velocities.
Fig. 13 shows that sturgeons swim around the aquarium under normal conditions while they swim up and down at a certain position in the aquarium under ammonia nitrogen stress [287]. Meanwhile, the phenomenon of sturgeon swimming around the aquarium under ammonia nitrogen stress has decreased. Key findings revealed significant changes in fish swimming patterns, spatial distribution, and velocity under ammonia stress, with species-specific responses to different concentrations. The method successfully differentiated normal behaviour from stress-induced abnormalities, demonstrating its potential for early detection of water quality issues in aquaculture settings. However, challenges such as turbid water and varying light conditions required additional solutions like adaptive image preprocessing techniques and regular maintenance protocols.
Figure 13: Behavioral trajectories of sturgeon under ammonia nitrogen stress. The five colours of the behavioural trajectory represent five fish (This figure was reproduced from [287]).
V Multimodal fusion in aquaculture
Multimodal data fusion has been used in aquaculture where data from various sources are integrated to understand fish behaviour, health, and environmental conditions. Recent advancements in other fields, such as deep learning, offer promising applications for aquaculture, particularly in fish tracking, counting, and behaviour analysis.
V-1 Multimodal fusion in fish tracking
For fish tracking, state-of-the-art vision-language models show significant potential. CLIP2Video [289] and AudioCLIP [290] extend the capabilities of their predecessors in image-text and audio-text matching by leveraging the contrastive language image pretraining (CLIP) model [291], and could be adapted for multimodal fish detection. For instance, the AudioCLIP architecture [290] could be modified to correlate visual fish features with sonar data, potentially improving detection accuracy in turbid waters commonly encountered in aquaculture settings. However, the significant differences between underwater RGB images and sonar images present unique challenges in multimodal tracking for aquaculture. Many existing multimodal tracking methods assume alignment between different modal spaces, using a prediction head to estimate the same target box across modalities. This assumption often fails in RGB-Sonar (RGB-S) tracking, where the representation of the target can vary significantly between the two imaging types [292, 293]. Furthermore, the spatial misalignment of multimodal features complicates the integration of information from these diverse sources. A critical challenge in multimodal fusion for aquaculture is data synchronization. Different sensors often have varying sampling rates and latencies. For instance, acoustic sensors might collect data at several kilohertz (kHz), while visual sensors typically capture frames at 30-60 Hz, and environmental sensors (like temperature or pH sensors) might only take measurements every few seconds or minutes.
Developing robust synchronization methods, such as timestamp alignment and interpolation techniques, is crucial for accurately fusing these diverse data streams in real-time aquaculture monitoring systems. To address these challenges, researchers have developed specialized approaches for underwater tracking. Li et al. [294] introduced a novel spatial cross-attention method that effectively bridges the gap between underwater RGB and sonar modalities. Their approach, which builds upon the Cross-Attention method [76], enables more nuanced interactive fusion between the different modalities. By accounting for the unique characteristics of each imaging type, this method achieves state-of-the-art performance in underwater tracking, demonstrating its potential for improving fish monitoring in complex aquaculture environments. The introduction of large-scale datasets like WebUOT-1M [295], which includes language prompts for video sequences, opens up new possibilities for underwater vision-language tracking. This could be particularly useful for developing more robust fish tracking systems in complex aquaculture environments, such as sea cages with varying light conditions and water clarity. However, a significant research gap remains in bridging vision-language models with sonar data for aquaculture applications, particularly in developing frameworks that can effectively leverage textual descriptions for enhanced underwater object detection and tracking. Future research should focus on developing more sophisticated multimodal fusion techniques to improve tracking accuracy and robustness in complex underwater environments.
V-2 Multimodal fusion in fish counting
In the realm of fish counting, recent developments in crowd counting using vision-language models offer promising approaches. CrowdCLIP [296] leverages pre-trained vision-language models to address the challenge of costly manual labelling, especially in dense scenes. This approach could be adapted for counting fish in crowded aquaculture tanks or ponds, where traditional methods often fail. CLIP-Count [297] presents an innovative end-to-end pipeline for estimating density maps of open-vocabulary objects with text guidance in a zero-shot manner. Its patch-text contrastive loss and hierarchical patch-text interaction module could be adapted to generate high-quality density maps for fish in various aquaculture scenarios, potentially improving counting accuracy in complex underwater environments with overlapping fish. Furthermore, VLCounter [298] explores the implicit association of semantic-patch embeddings of CLIP, showing improved performance in zero-shot object counting. This approach could be useful in aquaculture for counting different fish species or identifying specific fish features without extensive labelled datasets, a common challenge in diverse aquaculture operations. However, adapting these vision-language models to underwater environments presents unique challenges, particularly in handling water turbidity and varying lighting conditions. Future research should focus on developing specialized vision-language architectures that can better handle the specific challenges of underwater counting tasks while maintaining the benefits of zero-shot learning capabilities.
V-3 Multimodal fusion in fish behaviour analysis
In fish behaviour analysis, multi-sensor approaches combined with advanced algorithms offer new insights. The Multimodal Foundation Model (M3AE) [299], designed for general-purpose multimodal understanding, could be adapted to integrate visual, acoustic, and environmental data for comprehensive fish behaviour monitoring. This integration could enable early detection of stress or disease by correlating subtle behavioural changes with environmental parameters. Cui et al. [263] introduced a novel unified mixed-modality-based method for fish feeding behaviour assessment, capable of processing audio, visual, or audio-visual modalities. This approach demonstrates the potential of multimodal fusion to provide more robust and accurate assessments of fish behaviour than single-modality based methods. Similarly, Du et al. [300] showed that fusing information from audio, video, and imaging sonar modalities resulted in higher accuracy than using a single modality in assessing fish behaviour. While these studies demonstrate promising results, there remains a need to develop more sophisticated behavioural pattern recognition algorithms that can identify complex fish behaviours across different species and environmental conditions. Future research should explore integrating biological knowledge with multimodal deep learning approaches to better understand and interpret fish behavioural patterns in various aquaculture scenarios.
V-4 Challenges and solutions in multimodal fusion for aquaculture
Noise management is particularly challenging in underwater environments. Each modality (visual, acoustic) is subject to different types of noise. For example, visual data may be affected by turbidity and light scattering, while acoustic data can be distorted by ambient noise from equipment or other marine life. Advanced filtering techniques, such as Bernoulli filters [124] and differentiable particle filtering [123], need to be developed specifically for aquaculture applications to ensure reliable data fusion in these complex environments [301]. The relative importance and reliability of different data sources can vary based on environmental conditions and the specific behaviour being analyzed. Developing dynamic weighting schemes that adjust the contribution of each modality based on real-time quality assessments is essential [302]. For instance, acoustic data might be weighted more heavily in turbid conditions than visual data, while in clear water, the reverse might be true. Machine learning techniques, such as attention mechanisms in transformer-based models, offer promising approaches for adaptive weighting in multimodal fusion systems.
However, several challenges limit the direct application of these methods in aquaculture. Limited labelled data is a significant hurdle, as aquaculture-specific datasets are scarce compared to general computer vision tasks. Environmental variability in underwater conditions, such as changing turbidity, lighting, and temperature, can significantly impact sensor reliability and model performance. Additionally, species-specific behaviours require specialized models, as different fish species exhibit unique patterns [303]. The dynamic nature of fish movement and the three-dimensional space of aquatic environments further complicate the application of these technologies. Several approaches have shown promise in addressing these challenges. Few-shot learning techniques could help the models adapt to new species or environments with limited data [304], crucial for the diverse range of fish species in aquaculture. Self-supervised learning methods [305, 306] could leverage large amounts of unlabeled aquaculture data to pre-train robust models, potentially overcoming the scarcity of labelled datasets. Domain adaptation techniques [307] could help transfer knowledge from data-rich domains to aquaculture-specific tasks, bridging the gap between general computer vision tasks and the unique requirements of underwater environments.
Future research should focus on developing specialized datasets that capture the diversity of aquaculture environments and species. This could include multi-sensor data collected across different water conditions, fish life stages, and farming systems. Adapting advanced algorithms to underwater environments is crucial, particularly in handling variable lighting, water turbidity, and three-dimensional fish movement. Creating robust, generalizable models for diverse aquaculture scenarios will be key to the widespread adoption of these technologies. Integrating multiple data sources and advanced processing techniques can revolutionize fish monitoring, enhance production efficiency, and improve animal welfare in aquaculture operations. However, it is important to note that applying these technologies in real-world aquaculture settings may face practical challenges such as cost, maintenance of underwater sensors, and the need for real-time processing capabilities. As research progresses, critically evaluating the practicality and cost-effectiveness of these advanced multimodal approaches in various aquaculture settings will be essential.
VI Public dataset
@@
High-quality public datasets are crucial for developing and evaluating various methods including deep learning methods for fish detection, tracking, and behaviour analysis. Despite growing efforts, public datasets on underwater fish scenes are scarce, particularly in realistic aquaculture environments. This limitation has led many researchers to conduct analyses and behavioural studies under ideal or controlled conditions, which may not fully represent the challenges of real-world aquaculture settings. Table V summarizes the available public fish datasets, highlighting their diverse characteristics and limitations.
TABLE V: Summary of the various fish datasets.
@@
$2704\times 1520$ 86,452 Tracking [89]
FishTracker23
Approaching 1 million
boxes and frames
$1920\times 1080$
$1280\times 720$
850,000
Tracking [312]
WebUOT-1M
1,500 underwater
videos (total 10.5 hours)
and 408 categories
- 1,100,000 Tracking [295]
@@
FISHTRAC 14 videos $1920\times 1080$ 3,449
@@
The Fish4Knowledge dataset [308], captured in Taiwan waters from 2010 to 2013, offers a variety of marine scenes but suffers from low resolution (i.e. 320 × 240 pixels), limiting its applicability to current research needs. SeaCLEF2016 [309] builds on this with diverse species coverage and high-quality annotations, enhancing automated marine life identification capabilities. However, both datasets have limitations, including geographic specificity to Taiwanese waters, fixed camera positions that may miss certain behaviours, and low video quality issues due to underwater conditions. The DeepFish dataset [314] significantly contributes to underwater fish monitoring, containing 39,766 images from 20 locations in Australian marine environments. It offers annotations for multiple tasks including classification, counting, localization, and polygonal annotations for precise fish shape delineation. While being versatile, DeepFish initially lacks segmentation masks, limiting its use in instance segmentation tasks. Its diverse environmental coverage, from clear to turbid waters, enhances its utility for developing robust algorithms. However, the dataset’s focus on still images and specific marine ecosystems may limit its direct applicability to some aquaculture scenarios or dynamic behaviour studies.
The WebUOT-1M dataset [295] represents a significant advancement, offering a large-scale solution with 1.1 million frames across 1,500 underwater videos and 408 diverse target categories. Its inclusion of language prompts for each video facilitates multi-modal research and enables the development of more generalized and multi-modality underwater object-tracking models. However, the dataset’s web-sourced nature means it may include some biases in image quality, object representation, and environmental conditions. While this diversity can be beneficial for creating robust models, it may not always accurately represent real-world underwater scenarios encountered in specific research or industrial applications. For specialized research, datasets like 3DZeF20 [89] and AV-FFIA [263] offer unique perspectives. 3DZeF20 provides high-precision 3D data for zebrafish in laboratory settings, invaluable for detailed behaviour analysis but limited in ecological applicability. AV-FFIA innovatively combines hydrophone and video data for fish feeding behaviour analysis, offering deep insights into specific patterns. While all of them are invaluable for specific research questions, their applicability to broad aquaculture scenarios may be limited due to their focus on controlled environments or single species.
FishTrack23 [312] and FISHTRAC [105] represent the current state-of-the-art in real-world underwater object detection and tracking. These datasets address the complexities of natural marine environments, providing comprehensive annotations for multiple tasks. They excel in developing practical monitoring tools but may still face challenges with image quality consistency and global species representation. The BrackishMOT dataset[315] uniquely addresses tracking in turbid environments, which is crucial for many aquaculture settings. The Caltech Fish Counting (CFC) dataset [316] significantly contributes to underwater computer vision and marine ecology and is specifically designed to address the challenges of fish counting in complex underwater environments. Its diverse scenes capture different lighting conditions, water turbidity levels, and fish densities, making it an excellent resource for developing robust fish detection and counting algorithms. The dataset’s strength lies in its real-world applicability, presenting researchers with the genuine challenges faced in marine population surveys, such as occlusions, varying fish sizes, and species diversity. However, users should be aware that while the CFC dataset offers a comprehensive view of coral reef fish populations, it may have limitations in representing global marine ecosystems. The Fishsounds dataset [317], while offering audio data for various fish species, lacks sufficient data for comprehensive behaviour analysis. The Mullet Schools dataset [306] provides large-scale sonar imagery for fish counting but is species-specific, limiting its broad applicability in aquaculture research.
These datasets highlight several limitations in current public resources for aquaculture research. Many focus on specific conditions or species, not capturing the full range of aquaculture environments and diversity. Most are unimodal, lacking the integration of visual, acoustic, and environmental data crucial for comprehensive aquaculture monitoring. Varying annotation standards across datasets complicate cross-dataset evaluations. We suggest creating multi-modal datasets that integrate visual, acoustic, and environmental data specific to aquaculture settings to address these limitations and better serve future research needs. Existing datasets should be expanded to include a wider range of species and environmental conditions relevant to global aquaculture practices. Developing standardized annotation protocols would ensure consistency across aquaculture-focused datasets. Establishing benchmark datasets specifically designed for aquaculture applications, covering various farming systems and environmental conditions, would greatly benefit the field.
VII Challenges and future perspectives
@@
1) Massively available datasets: The wide application of intelligent technology in aquaculture, especially the success of deep learning algorithms in image processing [19], has highlighted the need for large labelled datasets. Although available datasets are gradually increasing, most are limited to identifying and detecting fish species. Open data on fish tracking, counting, and behaviour analysis is scarce. Passive acoustic monitoring is also gaining popularity for underwater listening [318, 255] and public sound data of underwater fish (e.g., Fishsound) are emerging. However, the sample size of these datasets has not yet reached critical mass. In the future, developing an international platform for sharing images and acoustic data will be essential to promote sustainable aquaculture development.
2) Edge computing and real-time processing: As aquaculture operations become more technologically advanced, there is an increasing need for real-time data processing and decision-making at the edge. Developing efficient algorithms and hardware solutions that can perform complex tasks like fish tracking, counting, and behaviour analysis on-site, without relying on cloud computing, will be crucial. This approach can reduce latency, improve data privacy, and enable faster responses to changing conditions in aquaculture environments. Future research should optimise existing algorithms for edge devices, develop specialized hardware for aquaculture applications, and create integrated systems that seamlessly combine multiple monitoring and analysis tasks in real-time.
@@
3) On-device machine learning: Most current fish tracking, counting, and behavioural analysis models are performed in the cloud or on high-performance GPUs. However, many aquaculture tasks require real-time responses, such as fish feeding and abnormal behaviour detection. Cloud-based models may struggle to guarantee this real-time performance, and many devices in remote and harsh aquaculture environments may not have consistent internet connectivity. On-device models can greatly reduce exercise pressure and make devices more intelligent, providing users with a better experience. However, terminal devices are often limited in processing power, power consumption, cost, and volume. Future developments could focus on reducing the complexity of computing and storage by optimizing neural network algorithms or compressing network models using techniques like knowledge distillation to enable their deployment on device chips.
@@
4) Integration of fish tracking, counting, and behaviour analysis: Most research addresses fish tracking, counting, and behaviour analysis as separate tasks. However, these tasks are often interconnected in real-world aquaculture scenarios and must be performed continuously in the same environment. Developing a joint model that can handle all three tasks simultaneously would be more memory-efficient and suitable for practical applications in aquaculture. A joint model would leverage the shared features and information among the tasks, reducing redundancy and improving overall performance. For example, accurate fish tracking can provide valuable counting and behaviour analysis information. In contrast, behaviour analysis can help identify and resolve tracking challenges such as occlusions and interactions between fish. More comprehensive and efficient systems could be developed for monitoring and managing aquaculture farms by integrating these tasks into a single framework. This approach would also reduce the computational resources required, making it more feasible to deploy such systems in real-world settings. Future research should focus on developing novel architectures and training strategies that can effectively combine fish tracking, counting, and behaviour analysis tasks.
@@
VIII Conclusions
This survey provides a comprehensive analysis of the current state of digital technologies in aquaculture, including vision-based sensors, acoustic-based sensors, and biosensors, for fish tracking, counting, and behaviour analysis. These technologies offer valuable tools for optimizing production efficiency, fish welfare, and resource management in aquaculture. However, each technology has its limitations, such as the sensitivity of vision-based sensors to environmental conditions, the high cost and complexity of acoustic-based sensors, and the potential invasiveness of biosensors. Despite the advancements in these technologies, significant challenges remain, including the scarcity of comprehensive fish datasets, the lack of unified evaluation standards, and the need for more robust and adaptable systems that can handle the complexities of real-world aquaculture environments. To address these challenges and drive progress in the field, future research should focus on developing diverse and representative datasets, establishing standardized evaluation frameworks, and exploring integrating multiple technologies to create more comprehensive and reliable monitoring systems. Emerging technologies such as multimodal data fusion, deep learning, and edge computing present exciting opportunities for advancing digital aquaculture. By leveraging these technologies, more accurate, efficient, and practical solutions can be developed for fish tracking, counting, and behaviour analysis, ultimately contributing to the sustainable growth and development of the aquaculture industry. As the field progresses, it is crucial to consider these technologies’ ethical implications and environmental impact. Developing solutions that are not only technologically advanced but also sustainable and respectful of animal welfare will be paramount for the future of aquaculture.
IX AUTHOR CONTRIBUTIONS